The cost of 5 boxes of envelopes and 5 packages of notebook paper is $25.38. How much would one box of envelopes cost?
1
Expert's answer
2019-02-04T08:53:05-0500

1. Let the cost of 1 box of envelopes be

"x"

. One package of notebook paper costs

"y"

. We see that

"5x+5y=25.38,""x+y=\frac{25.38}{5}=5.076."

2. One box of envelops hence costs

"x=5.076-y."

3. If you know the cost of one package of paper, simply substitute it for

"y"

. If you know the total cost of other quantity of boxes and packages, write an equation similar to the first one in pt.1, express

"y"

in terms of

"x"

and substitute it for

"y"

in the equation in part 2.